Given f(x) = 2x + 9, find ( f^-1 o f)(x)
Answer by HyperBrain(694)

f(x) = 2x + 9
Step 1: for convenience, let's replace f(x) with y
y=2x+9
Step 2: replace all y's with x and replace all x's with y
x=2y+9
Step 3: express x in terms of y
...
x=2y+9
2y+9=x
2y=x-9
y=(x-9)/2
Step 4: replace y with f-1(x)
f-1(x)=(x-9)/2
Power up,
HyperBrain!
My blogs:
pastime blogtime
hyperthesis
Go here